Grade 12th passTrigonometry

2tan-1(3/4)=??

alekhya
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
Write 2tan-1(3/4) =  tan-1(3/4) + tan-1(3/4) = k let

From formulae: 
tan-1(x + y) = tan-1(x) + tan-1(x)/(1 – tan-1(x)*tan-1(x)) 
Now put x = 3/4
k = 
tan-1(x + y)*(1 – tan-1(x)*tan-1(x)) =  tan-1(3/4 + 3/4)*(1 – tan-1(3/4)*tan-1(3/4)